# Weak Convexity of Fisher Information Matrix and Superresolved Localization of Blinking Sources of Light

### v2

@@ VI.3 Localization in one dimension
$\mathbf{Q}_{\text{blink}}=\mathbf{F}_{\text{blink}}=\frac{N}{2\sigma^{2}}\left(\begin{array}[]{cc}1+\delta&0\\ 0&1-\delta\end{array}\right).$ (72)
The quantum Fisher information matrix in the case of two cofluorescent incoherent sources with unknown relative brightness was found by Řehaček and coworkers in Ref. [28] for the same Gaussian shape of the impulse-response function. We assume that the relative brightness is known and use from this reference only the $2\times 2$ submatrix related to the centroid $S_{0}=(x_{1}+x_{2})/2$ and the separation $S=x_{1}-x_{2}$, which we denote by $\mathbf{Q}_{\text{coflu}}^{R}$. We transform this matrix into our parameters $x_{1}$ and $x_{2}$ as $\mathbf{Q}_{\text{coflu}}=A^{T}\mathbf{Q}_{\text{coflu}}^{R}A$, where $A$ is the transformation matrix for the (column) vectors of the parameters, $\{S_{0},S\}=A\{x_{1},x_{2}\}$, and obtain
$\mathbf{Q}_{\text{coflu}}=\frac{N}{2\sigma^{2}}\left(\begin{array}[]{cc}1+\delta-\beta&-\beta\\ -\beta&1-\delta-\beta\end{array}\right),$ (73)
@@
Figure 5: Variance lower bounds in the units of $L_{0}=2\sigma^{2}/N$ for estimating the unknown positions $(x_{1},x_{2})$ of two incoherent sources by measuring optimally $N$ photons in a microscope with the point-spread function of standard deviation $\sigma$. $\delta$ is the relative brightness difference of the sources. Parameters $(c,s)$ are the scaled centroid and separation, respectively. Parameters $(b,w)$ are the eigenparameters of the Fisher information matrix obtained from $(x_{1},x_{2})$ via a rotation by angle $\xi$.
The variance lower bounds $L_{ii}=(\mathbf{Q}^{-1})_{ii}$ are shown in Fig. 5(a) for the case of equal brightnesses $\mu_{1}=\mu_{2}=\frac{1}{2}$ and in Fig. 5(b) for the case of unequal brightnesses $\mu_{1}\neq\mu_{2}$. We see that, in contrast to Figs. 3(a) and (b), the variance lower bound in the cofluorescence scenario does not go to infinity at zero separation. This fact indicates the possibility of an efficient localization of two incoherent sources by means of optimal measurements, such as spatial mode demultiplexing [22, 46, 27, 47, 48]. In Fig. 5(a), the lower bound for the cofluorescent scenario lies above that for the blinking scenario, in correspondence with Eq. (66).
A similar behavior is shown in Figs. 5(c) and (d) for the scaled centroid and separation. The quantum Fisher information matrix is obtained in this basis as $\tilde{\mathbf{Q}}=O_{\pi/4}^{T}\mathbf{Q}O_{\pi/4}$, which gives the matrix $\tilde{\mathbf{Q}}_{\text{blink}}$ in the form of Eq. (49) and
$\tilde{\mathbf{Q}}_{\text{coflu}}=\frac{N}{2\sigma^{2}}\left(\begin{array}[]{cc}1-2\beta&-\delta\\ -\delta&1\end{array}\right).$ (75)
@@
Figure 6: Localization efficiency in the units of $H_{0}=N/2\sigma^{2}$ for estimating the unknown positions $(x_{1},x_{2})$ of two incoherent sources by optimally measuring $N$ photons in a microscope with the point-spread function of standard deviation $\sigma$ under condition of saturated quantum Cramér-Rao bound. The blinking scenario demonstrates a higher localization efficiency, whatever measure is used for the latter.
@@ VII Conclusion
Our comparison of cofluorescent and blinking scenarios for the task of localizing dim incoherent sources of light showed that the blinking scenario always has an informational advantage when the localization efficiency is measured by the average eigenparameter precision, which is a consequence of a fundamental property of the Fisher information matrix, its trace convexity proven in Theorem 1. This advantage is also proven for a more popular measure of efficiency, average position variance, however, only in the case of translation and rotation invariant microscopy and equal numbers of photons emitted by each source, which is a consequence of another fundamental property of Fisher information matrix, its weak convexity proven in the same theorem. Moreover, we have shown that this advantage is a consequence of the corresponding properties – trace and weak matrix convexity – of the quantum Fisher information matrix calculated for light on the image plane of a microscope.
We believe that these results are important for both the fields of superresolving optical microscopy and quantum metrology, at the junction of which they are located. On the one hand, they show that the information gain produced by blinking, well-known for decades in in the superresolving optical microscopy, is a mathematical consequence of such properties of the Fisher information matrix as additivity and convexity.
On the other hand, it shows that blinking of the sources can increase the Fisher information on their location even to a higher extent than the optimization of the quantum measurement, much studied in the last years in the framework of quantum metrology.

### v3

@@ VI.3 Localization in one dimension
$\underline{\mathbf{Q}}_{\text{blink}}=\underline{\mathbf{F}}_{\text{blink}}=\frac{N}{2\sigma^{2}}\left(\begin{array}[]{cc}1+\delta&0\\ 0&1-\delta\end{array}\right).$ (72)
The quantum Fisher information matrix in the case of two cofluorescent incoherent sources with unknown relative brightness was found by Řehaček and coworkers in Ref. [28] for the same Gaussian shape of the impulse-response function. We assume that the relative brightness is known and use from this reference only the $2\times 2$ submatrix related to the centroid $S_{0}=(x_{1}+x_{2})/2$ and the separation $S=x_{1}-x_{2}$, which we denote by $\underline{\mathbf{Q}}_{\text{coflu}}^{R}$. We transform this matrix into our parameters $x_{1}$ and $x_{2}$ as $\underline{\mathbf{Q}}_{\text{coflu}}=\mathbf{A}^{T}\underline{\mathbf{Q}}_{\text{coflu}}^{R}\mathbf{A}$, where $\mathbf{A}$ is the transformation matrix for the (column) vectors of the parameters, $\{S_{0},S\}=\mathbf{A}\{x_{1},x_{2}\}$, and obtain
$\underline{\mathbf{Q}}_{\text{coflu}}=\frac{N}{2\sigma^{2}}\left(\begin{array}[]{cc}1+\delta-\beta&-\beta\\ -\beta&1-\delta-\beta\end{array}\right),$ (73)
@@
Figure 6: Variance lower bounds in the units of $L_{0}=2\sigma^{2}/N$ for estimating the unknown positions $(x_{1},x_{2})$ of two incoherent sources by measuring optimally NN photons in a microscope with the point-spread function of standard deviation $\sigma$. $\delta$ is the relative brightness difference of the sources. Parameters $(c,s)$ are the scaled centroid and separation, respectively. Parameters $(b,w)$ are the eigenparameters of the Fisher information matrix obtained from $(x_{1},x_{2})$ via a rotation by angle $\xi$: $b=x_{1}\cos\xi+x_{2}\sin\xi$ and $w=x_{2}\cos\xi-x_{1}\sin\xi$.
The variance lower bounds $L_{ii}=(\underline{\mathbf{Q}}^{-1})_{ii}$ are shown in Fig. 6(a) for the case of equal brightnesses $\mu_{1}=\mu_{2}=\frac{1}{2}$ and in Fig. 6(b) for the case of unequal brightnesses $\mu_{1}\neq\mu_{2}$. We see that, in contrast to Figs. 4(a) and (b), the variance lower bound in the cofluorescence scenario does not go to infinity at zero separation. This fact indicates the possibility of an efficient localization of two incoherent sources by means of optimal measurements, such as spatial mode demultiplexing [22, 46, 27, 47, 48]. In Fig. 6(a), the lower bound for the cofluorescent scenario lies above that for the blinking scenario, in correspondence with Eq. (66).
Figure 7: Localization efficiency in the units of $H_{0}=N/2\sigma^{2}$ for estimating the unknown positions $(x_{1},x_{2})$ of two incoherent sources by optimally measuring NN photons in a microscope with the point-spread function of standard deviation $\sigma$ under condition of saturated quantum Cramér-Rao bound. The blinking scenario demonstrates a higher localization efficiency, whatever measure is used for the latter.
A similar behavior is shown in Figs. 6(c) and (d) for the scaled centroid and separation. The quantum Fisher information matrix is obtained in this basis as $\tilde{\underline{\mathbf{Q}}}=\mathbf{O}_{\pi/4}^{T}\underline{\mathbf{Q}}\mathbf{O}_{\pi/4}$, which gives the matrix $\tilde{\underline{\mathbf{Q}}}_{\text{blink}}$ in the form of Eq. (49) and
$\tilde{\underline{\mathbf{Q}}}_{\text{coflu}}=\frac{N}{2\sigma^{2}}\left(\begin{array}[]{cc}1-2\beta&-\delta\\ -\delta&1\end{array}\right).$ (75)
@@ VII Conclusion
Our comparison of cofluorescent and blinking scenarios for the task of localizing dim incoherent sources of light showed that the blinking scenario always has an informational advantage when the localization efficiency is measured by the average eigenparameter precision, which is a consequence of a fundamental property of the Fisher information matrix, its trace convexity proven in Theorem 1. This advantage is also proven for a more popular measure of efficiency, average position variance, however, only in the case of translation and rotation invariant microscopy and equal numbers of photons emitted by each source, which is a consequence of another fundamental property of Fisher information matrix, its weak convexity proven in the same theorem. Moreover, we have shown that this advantage is a consequence of the corresponding properties – trace and weak matrix convexity – of the quantum Fisher information matrix calculated for light on the image plane of a microscope. The informational advantage of the blinking scenario can be ascribed to a higher a priori information, which one has about the number of simultaneously active emitters: 1 in the blinking scenario and $K>1$ in the cofluorescent one. The advantage in a priori information can be quantified by the advantage in the Fisher information calculated in this work.
We believe that these results are important for both the fields of superresolving optical microscopy and quantum metrology, at the junction of which they are located. On the one hand, they show that the information gain produced by blinking, well-known for decades in the superresolving optical microscopy, is a mathematical consequence of such properties of the Fisher information matrix as additivity and convexity. On the other hand, it shows that blinking of the sources can increase the Fisher information on their location even to a higher extent than the optimization of the quantum measurement, much studied in the last years in the framework of quantum metrology.